OpenStudy (anonymous):

Y=5-2x+3x standard form

OpenStudy (anonymous):

y = mx +c is the standard form ...where Y and X are variable and c is constant now solve the given equation in that form keeping Y on one side and X and C on other side .

OpenStudy (joannablackwelder):

I think standard for is ax + by = c.

OpenStudy (joannablackwelder):

So, you would need to combine the like terms and move the x's and y's to the same side of the equals sign.

OpenStudy (the_fizicx99):

@vinurules you're wrong its Ax + By = C Y=5-2x+3x Y = 5 - 1X 1X +1 So its, x + y = 5

OpenStudy (anonymous):

@JoannaBlackwelder -depeneds -- y = mx + c is a standard equation for a straight line,ie one degree equation so it can be used. so i solve it as Y = 5 -2X + 3X Y = 5 - X (2-3 ) Y = 5 -X (-1) Y= 5 +X Y = X +5 OR even u ur form it is y-x = 5 or x -y = -5 both are same... ( sorry @tHe_FiZiCx99 made a calculation mistake i belive )

OpenStudy (the_fizicx99):

Oh wait typo, Y = 5 + 1X You would subtract to move x to the left, though you can't have a negative so you multiply it by (-1) to cancel it out. Y = 5 + 1X -1x -1x -1x + y = 5 -1x(-1) + y(-1) = 5(-1) x - y = -5

OpenStudy (the_fizicx99):

@vinurules No, I made a typo.

OpenStudy (anonymous):

@tHe_FiZiCx99 -same way mistake is a mistake..i feel. rest i will by what way u get marks by your teacher is a right way.. but y = mx +c m = slope of the line and c = value of intercept is a easiest to graph it .

OpenStudy (the_fizicx99):

@vinurules Yes, I made a mistake, I clicked the wrong sign... It's normally used y = mx + b though b and can be any variable. It'll still be the y intercept..

OpenStudy (anonymous):

@tHe_FiZiCx99 - in your equation , b can never be variable .it will be constant value only for a staright line eqution .

OpenStudy (the_fizicx99):

Yes, it can also be C, it'll still be the y-intercept.

OpenStudy (the_fizicx99):

I meant to say you can call it any letter in the alphabet. Sorry.